OpenStudy (anonymous):

15-3d/2d-10 (simplify the expression)

OpenStudy (anonymous):

15-3d 3-3d -------= ------- 2d-10 2d-2

OpenStudy (anonymous):

3( 5-d) -------- = -1 1/2 or -1.5 -2(5-d)

OpenStudy (anonymous):

@gitahimart it would be -3(d-5)

OpenStudy (anonymous):

it can also be so, this will not alter the outcome, check keenly: -3(d-5) ------- = -3/2 = -1 1/2 or -1.5 2 (d-5)

OpenStudy (anonymous):

You're right :)

OpenStudy (anonymous):

:) you're right too